OpenStudy (anonymous):

find the intgerating factor fot first linear differential equation dy/dt +(1+t)*exp(-t)*y = 0

OpenStudy (irishboy123):

its \(exp [ \int 1+t).e^{-t} dt]\) but you can separate this too

OpenStudy (astrophysics):

\[y'+(1+t)e^{-t}y=0\] notice it's in the form \[y'+a(x)y=b(x)\] where b(x) = 0 so the integrating factor for linear equations is \[\huge p(x) = e^{\int\limits a(x) dx}\]

OpenStudy (anonymous):

this my solution \[ e ^{-2e ^{-t}-te ^{-t}} \] is this the last solution or is there still any step that can be done

OpenStudy (astrophysics):

Looks good

OpenStudy (astrophysics):

That is just your integrating factor, you have not however found the general solution for your differential equation

OpenStudy (mathmale):

Better do a quick review of "integrating factors," including where they come from and how they are used in solving differential equations.

OpenStudy (michele_laino):

the general solution of this ODE: \[\Large u' + P\left( t \right)u = Q\left( t \right)\] is the subsequent function: \[\large \begin{gathered} u\left( t \right) = \exp \left( {\int { - P\left( t \right)dt} } \right)\left\{ {k + \int {Q\left( t \right) \cdot \exp \left( {\int {P\left( t \right)dt} } \right)} } \right\} \hfill \\ \hfill \\ k \in \mathbb{R} \hfill \\ \end{gathered} \] please see my tutorial: http://openstudy.com/study#/updates/552b4f50e4b04e5707c1d380 now, in our case, we have: \[\Large Q\left( t \right) \equiv 0\] so, we get, as solution: \[\Large \begin{gathered} u\left( t \right) = k\exp \left( {\int { - P\left( t \right)dt} } \right) \hfill \\ k \in \mathbb{R} \hfill \\ \end{gathered} \] furthermore, a simple integration gives the subsequent primitive function: \[\Large \int { - P\left( t \right)dt} = {e^{ - t}}\left( {t + 2} \right)\]